# Pulsation of Burner-Stabilized CH4-O2 Flames Moderated by CO2 Addition – Frequencies, Modes and Regime Diagrams

### v3

@@ 3.3 Multi-mode oscillation
Two more complex cases of pulsating flames are presented in Fig. 7, including: (a) single-mode oscillations coupled with flickering instabilities, and (b) multi-mode oscillations. In the former case, the dominant flame oscillation mode remains a centripetal traveling wave, but flickering instabilities introduce a slow edge-mode perturbation that modulates its relative phase without significantly affecting its amplitude and spatial distribution. In the latter case, the amplitude of the flame oscillation varies significantly over time.
For example, in a representative measurement shown in Fig. 7(b), the integrated chemiluminescence intensity exhibits a clear beat pattern, indicating the presence of two fundamental oscillation frequencies interfering with each other. During the time interval of constructive interference, the spatial pattern of the flame oscillation remains axisymmetric, while the propagation of the heat release rate perturbation becomes bidirectional – waves originate from a ring zone and travel toward both the center and the edge of the burner. In contrast, during the time interval of destructive interference, the flame oscillation pattern appears drastically different: perturbation waves travel along two separate loops near the top and bottom edges of the burner and meet along the diameter. More complex interactions among three or more modes have also been observed at much higher heat release rates.
@@ 3.4 Regime diagrams
The primary frequencies and modes of pulsating CH4-O2-CO2 flames, measured under various fuel flow rates, equivalence ratios, and dilution factors, are provided in Supplementary Material I. These results are also summarized in regime diagrams and frequency contour plots, as displayed in Figs. 8 and 9. Under the current experimental configuration, pulsating flame instability was observed only under fuel-rich conditions, specifically within the equivalence ratio range of 1.18 to 2.50. The range of unstable equivalence ratios narrows with increasing CO2 dilution (decreasing $\eta$), while its dependence on the CH4 flow rate is non-monotonic – first expanding with increasing CH4 flow rate, then peaking around 0.4 SLPM, and finally shrinking to a single point. This trend is consistent with the results reported previously in [27]. The highest CH4 flow rate at which flame pulsation was observed occurs near 1.32 SLPM, with the corresponding equivalence ratio and dilution factor being $\phi$ = 1.48 and $\eta$ = 1.00, respectively.
Figure 8: Regime diagrams of instability modes at selected CH4 mass flow rates. Lines indicate the observed stability boundaries.
@@
[…]
@@ 4 Conclusions
The pulsating instabilities of premixed CH4-O2 flames on a porous-plug burner were investigated at various levels of CO2 dilution, across a wide range of equivalence ratios, and flow rates. Time-resolved measurements of flame chemiluminescence and average temperature were obtained, from which the primary oscillation frequencies of the pulsating flames were determined using fast Fourier transform and harmonic power analysis. Additionally, spatiotemporally resolved measurements of gas temperature were conducted using spectrally modulated laser absorption diagnostics synchronized with high-speed chemiluminescence imaging.
From the results of the current measurements, a mode transition phenomenon in the pulsating flames was revealed through phase-locked analysis. Under fuel-rich conditions with relatively low heat release rates and flow rates, the flames exhibited quasi-periodic pulsation in a single mode. At elevated flow rates, flickering instability appeared and began to interfere with the thermo-diffusive pulsation, creating sidebands around the primary frequency. As the heat release rate increased, mode splitting occurred, and the flames eventually pulsated in multiple modes. The stability boundaries and regime diagrams of the flame oscillation modes were obtained, and representative examples of different modes were analyzed.
Although obtained from experiments on a specific porous-plug burner, the results of the current study – such as the overall trends of the stability boundary, primary frequency, and mode distributions – are expected to be broadly applicable to other CO2-diluted oxy-fuel combustion systems. They provide a basis for theoretical and numerical extrapolation toward applications relevant to practical combustion devices, where flame stability is a critical design consideration. While the current study primarily focused on methane flames, the underlying concepts can be easily extended to other fuels such as hydrogen and ammonia. In this sense, the present study should prove valuable for both fundamental research on flame dynamics and practical applications of CO2-moderated oxy-combustion.
@@ Acknowledgments
This research is supported by the National Natural Science Foundation of China under Grants No. 92152108 and No. 12472278, and the Space Application System of China Manned Space Program under projects ”Nonlinear Dynamics of Flame Instability under Microgravity Conditions” and ”Ignition Mechanism of Premixed Near-Limit Flames under Microgravity Conditions”.

### v4

@@ 3.3 Multi-mode oscillation
Fig. 7 shows two representative cases of pulsating flames exhibiting more complex dynamics, including: (a) single-mode oscillations coupled with flickering instabilities, and (b) multi-mode oscillations. In the former case, the dominant flame oscillation mode remains a ring-shaped traveling wave, but flickering instabilities introduce a slow edge-mode perturbation that modulates its relative phase without significantly affecting its amplitude and spatial distribution. In the latter case, the amplitude of the flame oscillation varies more drastically over time.
For example, in a representative measurement shown in Fig. 7(b), the integrated chemiluminescence intensity exhibits a clear beat pattern, indicating the presence of two fundamental oscillation frequencies interfering with each other. During their constructive interference, the spatial pattern of the flame oscillation appears axisymmetric – perturbation waves emerge from a ring zone and travel toward both the center and the edge of the burner. In contrast, during their destructive interference, the flame oscillation pattern appears remarkably different: perturbation waves travel along two separate loops near the top and bottom edges of the burner and converge along the diameter. More complex interactions among three or more modes have also appeared at conditions of higher heat release rates.
In the present study, a flame was classified as stable, flickering, single-mode, or multi-mode pulsating based on three parameters: (1) the dominant/strongest frequency, $f_{osc}$, defined by Eqn.3; (2) the signal-to-noise ratio (SNR) at $f_{osc}$, evaluated in relative units of dB with respect to the average noise floor ($\hat{S}_{noise}$ = -80 dB/Hz) across 100 - 500,000 Hz, i.e., SNR $\hat{S}(f_{osc})-\hat{S}_{noise}$; and (3) the relative spectral power ($R$) of the second-strongest pulsating frequency ($f_{osc,2}$) with respect to the strongest frequency ($f_{osc}$), i.e., $R=\hat{S}(f_{osc,2})-\hat{S}(f_{osc})$. First, a flame was classified as oscillating if SNR >> 30 dB; otherwise, it was considered stable. It is worth noting that the main conclusions of the present study were not sensitive to the specific choice of this threshold, as the qualitative results remained largely unchanged when its value was varied within $\pm$ 10 dB. Then, an oscillating flame was classified as flickering if $f_{osc}<$ 20 Hz; otherwise, it was classified as pulsating. As illustrated in Fig. 8, this classification between the flickering mode and the pulsating mode is self-explanatory. Finally, a pulsating flame was classified as multi-mode if $R>$ -10 dB (representing that the spectral power of the second strongest mode is higher than 10% of the strongest mode); otherwise, it was classified as single-mode. Note that, in principle, the transition between single-mode and multi-mode should be continuous and does not necessarily involve a definite boundary (in a sense, the value of $R$ itself can be regarded as an objective metric of mode purity). However, to facilitate discussion of the current results, a -10 dB criterion was imposed. As demonstrated later in Fig. 9, this threshold value closely captures the steepest slope of $R$ with respect to $\phi$, representing a reasonably good separation of modes.
[…]
Figure 8: Joint distribution of the observed $f_{osc}$ and SNR from the current experiments.
@@ 3.4 Regime diagrams
The primary frequencies and modes of pulsating CH4-O2-CO2 flames, measured under various fuel flow rates, equivalence ratios, and dilution factors, are provided in Supplementary Material I. These results are also summarized in regime diagrams and frequency contour plots, as displayed in Figs. 9 and 10. Under the current experimental configuration, pulsating flame instability was observed only under fuel-rich conditions, specifically within the equivalence ratio range of 1.18 to 2.50. The range of unstable equivalence ratios narrows with increasing CO2 dilution (decreasing $\eta$), while its dependence on the CH4 flow rate is non-monotonic – first expanding with increasing CH4 flow rate, then peaking around 0.4 SLPM, and finally shrinking to a single point. This trend is consistent with the results reported previously in [26]. The highest CH4 flow rate at which flame pulsation was observed occurs near 1.32 SLPM, with the corresponding equivalence ratio and dilution factor being $\phi$ = 1.48 and $\eta$ = 1.00, respectively.
Figure 9: Regime diagrams of instability modes at selected CH4 mass flow rates. Solid lines: the observed stability boundaries. Dash lines: the boundary between single-mode and multi-mode oscillations defined by R=10R=-10 dB.
@@
The dependence of the effective Lewis number $Le$ on the dilution factor $\eta$ and the equivalence ratio $\phi$ was evaluated from 1-D freely-propagating adiabatic flame simulations performed with Cantera [15] using the DRM19 mechanism [23]. The results are shown in Fig. 11. Across the range of conditions explored in the current study, $Le$ varies between 0.96 and 1.09.
Figure 11: Effective Lewis number ($Le$) as a function of equivalence ratio ($\phi$) and dilution factor ($\eta$), calculated using the DRM19 mechanism [23].
A non-monotonic dependence of $Le$ on CO2 addition is observed, arising from competing effects of mass and thermal diffusion. On the one hand, CO2 addition reduces the mixture-averaged species diffusivity and slows mass transport; on the other hand, it increases the mixture’s specific heat capacity and reduces the overall thermal diffusivity, especially at high dilution. The reduction in mass diffusivity dominates at low CO2 dilution ($0.8\leq\eta\leq 1.0$), producing a negative dependence of $Le$ on $\eta$; whereas at relatively high CO2 dilution ($\eta<0.8$), the large CO2 heat capacity becomes the dominant factor and yields a positive dependence of $Le$ on $\eta$. At conditions of relatively high $Le$ (e.g., $Le>1.08$), the flame pulsation was multi-mode. Throughout the current study, no flame pulsation was observed at conditions of $Le<1$.
@@ 4 Conclusions
The pulsating instabilities of premixed CH4-O2 flames on a porous-plug burner were investigated at various levels of CO2 dilution, equivalence ratios, and flow rates. Heat loss from the flat burner surface amplified the thermo-diffusive instability of these flames and lead to pulsations at near-unity Lewis numbers where freely propagating flames would be stable.
In the present study, different modes of flame pulsation were observed. Under fuel-rich conditions with relatively low heat release and flow rates, the flames exhibited quasi-periodic pulsation in a single mode. At elevated flow rates, buoyancy-controlled, low-frequency flickering instability appeared and began to interfere with the thermo-diffusive pulsation, creating sidebands around the primary frequency. At elevated heat release rates, nonlinear interaction between flame flicker and single-mode pulsation further triggered mode splitting, and multi-mode pulsation emerged. The stability boundaries and regime diagrams of the flame oscillation modes were established, and representative measurements of the spatiotemporal variations in pulsating flames were analyzed.
The general trends of the stability boundary and mode distributions are generalizable to other CO2-diluted oxy-fuel combustion systems. While this study focuses on methane flames, the underlying concepts can be extended to other fuels, such as hydrogen and hydrogen-blended fuels. In this sense, the present study should prove valuable for both fundamental research on flame dynamics and practical applications of oxy-combustion.
@@ Acknowledgments
This research is supported by the National Key Research and Development Program of China under Grant No. 2025YFF0511801, the National Natural Science Foundation of China under Grants No. 12472278 and No. 92152108, the Space Application System of China Manned Space Program under projects ”Nonlinear Dynamics of Flame Instability under Microgravity Conditions” and ”Ignition Mechanism of Premixed Near-Limit Flames under Microgravity Conditions”, and the Open Research Program of National Key Laboratory of Fundamental Algorithms and Models for Engineering Simulation.